Oxygen gas with a density of 3kgm-3 exert a pressure of 3*106 Pa on the walls of its container. What is the root mean square velocity of oxygen molecules?

 

GK


1
Expert's answer
2021-09-06T12:42:39-0400

We know that: "P=\frac{1}{3}\rho v^2" where "P" is pressure, "\rho" is density, "v" is velocity.

"3\times10^6=\frac{1}{3}\times3\times v^2\\
\Rightarrow v^2=3\times 10^6\\
\Rightarrow v=\sqrt3\times10^3 m/s\\
\Rightarrow v=1000\sqrt3 \ m/s"